OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.) help me my teacher reassign this one

OpenStudy (anonymous):

i wish i knew what "one segment" meant

OpenStudy (anonymous):

Saitleite i forgot the answer.

OpenStudy (anonymous):

diameter of the circle is the diagonal of the square. since the square has side 6, we know the diameter of the circle is \[d=6\sqrt{2}\] and so the radius is \[3\sqrt{2}\] area of square is 36 area of circle is \[\pi r^2=\pi (3\sqrt{2})^2=18\pi\]

OpenStudy (anonymous):

the diameter of circle is6 sqrt(2) and radius is diameter /2

OpenStudy (anonymous):

so we know the area of the square, and we know the area of the circle. if we knew what 'ONE SEGMENT' meant i we would be done

OpenStudy (anonymous):

THE ANSWER TO THE QUESTION!

OpenStudy (anonymous):

oooooooooooooh maybe it means one of those four parts? not sure, but if it does then answer is \[\frac{18\pi-36}{4}\]

OpenStudy (anonymous):

or if you prefer \[\frac{18(\pi-2)}{4}=\frac{9(\pi-2)}{2}\] but a picture would really help here.